To study the causes of an outbreak of aflatoxin poisoning in Africa, investigators conducted a case-control study with 40 case-patients and 80 controls. Among the 40 poisoning victims, 32 reported storing their maize inside rather than outside. Among the 80 controls, 20 stored their maize inside.

a. Set up a 2 x 2 table and calculate the odds ratio for the association between inside storage of maize and illness.

b. Interpret the odds ratio.
Answer #1

Case control study definition

A study that compares patients who have a disease or outcome of interest (cases) with patients who do not have the disease or outcome (controls), and looks back retrospectively to compare how frequently the exposure to a risk factor is present in each group to determine the relationship between the risk factor and the disease.

A) Set up a 2 x 2 table and calculate the odds ratio for the association between inside storage of maize and illness.

Sl.no Cases (ill) Controls(not ill) Total
Exposed( Maize stored Inside) 32 (A) 20 (B) A+B=52
Unexposed( Maize Stored outside) 8 (C) 60 (D) C+D=68
Total A+C= 40 B+D=80 A+B+C+D=120

Odd ratio = (A*D) / (B*C)

= (32*60) / ( 20*8)

= 1920 / 160

= 12

...................................................................................................................................................................

B) Odds ratio

Definition

The Odds Ratio is a measure of association which compares the odds of disease of those exposed to the odds of disease those unexposed.

Interpretation of odds ratio

  • OR of 1 would suggests that there is no difference between the groups; i.e. there would be no association between the suggested exposure (fish) and the outcome (being ill)
  • OR of > 1 suggests that the odds of exposure are positively associated with the adverse outcome compared to the odds of not being exposed
  • OR of < 1 suggests that the odds of exposure are negatively associated with the adverse outcomes compared to the odds of not being exposed. Potentially, there could be a protective effect

In the example above, we can conclude those who stored maize inside( exposed) were 12 times ( OR=12) more likely to aflatoxin poisoning compared to those who stored maize outside ( unexposed ).